OpenStudy (anonymous):

evaluate the integral. S from 3 to 0; (2x^2 + x + 3)dx

OpenStudy (anonymous):

\[[(2/3)x ^{3} +(1/2)x ^{2} +3x ] \] between limits 3..0 so [ ( 2/3) (27) +(1/2)(9) +3(3) ] - [0] = 18 + 9/2 +9 = 31.5